OpenStudy (xapproachesinfinity):

given the equation \[f(x+y)=f(x)+f(y)+x^2y+xy^2\] and the limit: \[\lim_{x \to 0} \frac{f(x)}{x}=1\] a) find f(0) b) find f'(0) c) find f'(x)

OpenStudy (xapproachesinfinity):

from the limit i concluded that f'(0)=0 and than happen if f(0)=0 considering that limit to be the f'(0)=lim (f(x)-f(0))/x

OpenStudy (xapproachesinfinity):

is what i did sound fine?

OpenStudy (xapproachesinfinity):

@Kainui

OpenStudy (xapproachesinfinity):

i meant f'(0)=1 not 0

OpenStudy (xapproachesinfinity):

now i'm thinking to let y=x to get \[f(2x)=2f(x)+2x^3\] doing the derivative \[2f'(2x)=2f'(x)+6x^2\] \[f'(2x)=f'(x)+3x^2\] not sure where this leads to eheh

OpenStudy (xapproachesinfinity):

hmm i think i didn't get that f(0) correct!

OpenStudy (xapproachesinfinity):

thinking...

ganeshie8 (ganeshie8):

this is @freckles 's fav problem!

OpenStudy (xapproachesinfinity):

hehe where is @freckles

OpenStudy (xapproachesinfinity):

i will come back later to think about this just leave me here your notes :)

OpenStudy (freckles):

\[f(x+y)=f(x)+f(y)+x^2y+xy^2\] \[\text{ put } x=y \] \[f(x+x)=f(x)+f(x)+x^3+x^3\] \[f(2x)=2f(x)+2x^3\] \[f(2(0))=2f(0)+2(0)^3\] \[f(0)=2f(0)+0\] \[f(0)-2f(0)=0\] \[-1f(0)=0\] \[f(0)=0\] But I bet you were trying to skip to the derivative part \[\lim_{x \rightarrow 0}\frac{f(x)-f(0)}{x-0}=(f(x))'|_{x=0} \\ \text{ and we have } f(0)=0 \\ \text{ so we have } \lim_{x \rightarrow 0} \frac{f(x)}{x}=(f(x))'|_{x=0}\] But we were given: \[\lim_{x \rightarrow 0} \frac{f(x)}{x}=1\] so f'(0)=1 ... now I have to think about the f'(x) again

OpenStudy (freckles):

http://openstudy.com/users/myininaya#/updates/5410d734e4b08b11f1302400 check out what @ganeshie8 wrote towards the end :)

OpenStudy (xapproachesinfinity):

Interesting I used that differently I found f'(x) but I'm in the phone now I can't post it once I get home I will post it

OpenStudy (xapproachesinfinity):

Eh that's exactly the same question lol Anyways I will post my f'(x) later

OpenStudy (freckles):

and i think you posted in that question lol

ganeshie8 (ganeshie8):

Ahh I remember now xD @xapproaches , here is a copy paste from that thread for computing \(f'(x)\) : \[f(x+y)=f(x)+f(y)+x^2y+xy^2\] \[f(x+y)-f(x) = f(y)+x^2y+xy^2\] divide through out by \(y\) \[\dfrac{f(x+y)-f(x)}{y} = \dfrac{f(y)}{y}+x^2+xy\] take the limit and \(y\to 0\) and stare at left hand side \[\lim\limits_{y\to 0}~\dfrac{f(x+y)-f(x)}{y} = \lim\limits_{y\to 0} \left[ \dfrac{f(y)}{y}+x^2+xy\right]\] \[f'(x) = 1 + x^2 + 0\]

OpenStudy (xapproachesinfinity):

eh i didn't remember actually i was just trying some problems from stewart textbook! here is what i did \[\large f'(x)=\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}=\lim_{h\to0}\frac{f(x)+f(h)+x^2h+xh^2-f(x)}{h}\] \[\large =\lim_{h\to0}\frac{f(h)+x^2h+xh^2}{h}=\lim_{h\to0}[\frac{f(h)}{h}+x^2+hx]\] \[\large =1+x^2\] which the same result

OpenStudy (xapproachesinfinity):

i don't why i didn't see that f(0) even thought i did take y=x hehe

OpenStudy (xapproachesinfinity):

though*

OpenStudy (freckles):

mad skills :p @xapproachesinfinity

OpenStudy (xapproachesinfinity):

damn i feel down that i didn't see that f(0) hehe

OpenStudy (kainui):

Woah, for some reason that method is really quite awesome to use (x+y) and make it (x+h), it just feels so nice.... XD

OpenStudy (xapproachesinfinity):

:) i will post a different question and how can come up with the solution

OpenStudy (kainui):

I have one question for you, what is f(x) itself? (This isn't too far away I think)

OpenStudy (xapproachesinfinity):

well just take integral yes?

OpenStudy (xapproachesinfinity):

we have f(0)=0 already so we can find C

OpenStudy (xapproachesinfinity):

c=-1 hmm

OpenStudy (xapproachesinfinity):

oops no le me see

OpenStudy (xapproachesinfinity):

c=0 yes

OpenStudy (xapproachesinfinity):

okay let me post it \[f(x)=x+\frac{ x^3}{3}\]

OpenStudy (kainui):

Yeah, =P

OpenStudy (xapproachesinfinity):

:)

OpenStudy (xapproachesinfinity):

somehow you like integrals hehe :P

OpenStudy (kainui):

Integrals are probably the coolest thing I've ever seen in my entire life. Learning solids of revolution was pretty much a spiritual awakening for me, and that is not even (or oddly) an exaggeration. I mean, a single simple integral allows you to take a circle, imagine that it has an infinite number of smaller rings inside of it. But then you actually get to add up all those circles of "0" width inside together, multiplied by their circumferences (just baby's length*width) to get the area of the circle. Then people come along and say they invented with axioms, what a bunch of cheap liars. Not only that, but they're too afraid to say the truth that division by zero is defined, it's just such a sensitive subject we have to call it "calculus" as if it's separate... Ok now I'm rambling but yeah, somehow I like integrals and no one else seems to realize how incredible they truly are! haha.

OpenStudy (xapproachesinfinity):

haha, that's deep ^^

OpenStudy (xapproachesinfinity):

why do you call them cheap though, didn't they do a lot of awesome stuff hehe

OpenStudy (kainui):

Depends on who we're talking about lol

OpenStudy (xapproachesinfinity):

idk who you were talking about heheh, at any rate must be some mathematicians lol

OpenStudy (kainui):

I mean I'm not saying axioms can't be helpful, I just think people believe that math comes from axioms. It's useful for creating a giant system based on pure logic so that you can compare it to similar mathematical structures. But I don't believe the set theory construction of calculus is actually calculus. I'm not really talking about anyone in particular anyways haha, like everyone has their own way and if they like/need/want/use axioms let them. I just don't haha.

OpenStudy (freckles):

so you believe 5/0 is defined if so then how do you define it?

OpenStudy (kainui):

I never said that.

OpenStudy (freckles):

" Not only that, but they're too afraid to say the truth that division by zero is defined, " oh you were talking about 0/0

OpenStudy (freckles):

maybe

OpenStudy (xapproachesinfinity):

well i must admit axioms are not so convincing sometimes or perhaps i just don't got it hehe

OpenStudy (xapproachesinfinity):

well if you define 0/0 then you need to define 5/0 as well?

OpenStudy (freckles):

not really 0/0 is indeterminate form some cases you can define it and some cases you can't

OpenStudy (freckles):

I think anyways

OpenStudy (kainui):

Yeah, I am not saying division by zero is defined for all values, just stuff like 0/0. Basically all of calculus is just dividing 0 by 0 and doing stuff that's scary. It depends on what the context is, does what you're doing with the division by zero have some meaning behind it or is it worthless nonsense?

OpenStudy (freckles):

ok well that was the only thing that scared me of your whole paragraph

OpenStudy (kainui):

I don't see any reason why things have to have definitions at all in mathematics. If you do something and what you get out of it is a useful result and the interpretation you get out of it gives you something else useful, why fight it? I am just not interested in rigor at all, and usually proofs are not interesting to me because of several really stupid reasons that I will spare you lol.

OpenStudy (freckles):

Are you a mathematician or an engineer?

OpenStudy (freckles):

or i mean I guess you can be both

OpenStudy (kainui):

Neither haha.

OpenStudy (freckles):

like i met an engineer they say they didn't have to care about proofs just results and you remind me of her

OpenStudy (kainui):

Well there's a difference between why an engineer doesn't like proofs and why I don't like proofs.

OpenStudy (xapproachesinfinity):

hehe math with no proofs is just not working no?

OpenStudy (kainui):

I am kind of being misunderstood here when I say these things, but I'm not sure I am really interested in making the point I'm trying to say here haha. Let's just say I like "good" proofs.

ganeshie8 (ganeshie8):

Kainui real analysis is all about rigor. It starts with construction of rational numbers and builds all the way up to integrals. It is more like language of mathematics, defining each term and proving every little thing along the way. I have fallen in love with real analysis because of its rigor despite my 10 year brain damaging engineering experience lol

ganeshie8 (ganeshie8):

you're talking about real analysis when you talk about quality of proofs

OpenStudy (kainui):

Yeah but isn't real analysis kind of fake compared to complex analysis? I just can't get into it, I feel like there's kind of a lot of broken arguments going on in real analysis since they are depriving themselves of a fuller picture. Where do axioms come from? They are also kind of redundant, which bothers me, it just doesn't feel right or like real mathematics at all. All the definitions in analysis seem just arbitrary to me and not natural in any way, it just makes me too uncomfortable to study. When I think back to when I took real analysis we were told to prove things like that there are an infinite number of integers, and use epsilon delta proofs which seems to just tuck infinities away rather than actually dealing with them. It just feels like a big meaningless game to me.

ganeshie8 (ganeshie8):

I like both, but i guess i like real analysis more as it is my first love xD

OpenStudy (xapproachesinfinity):

you guys are mathematicians :)

OpenStudy (kainui):

In my opinion you are what you do everyday. And I do math every single day, so in that sense I am a mathematician I'll agree haha. @xapproachesinfinity

OpenStudy (xapproachesinfinity):

hehe everyday without exceptions yeah i would agree no less if you live with math like that